OpenStudy (inowalst):

The quadratic expressions \(x^2+5x−14\) and \(49-x^2\)​​ share a common binomial factor.

OpenStudy (yanasidlinskiy):

So, what's the question? Are we supposed to factor it out?

Parth (parthkohli):

They do.

OpenStudy (inowalst):

Yes. @YanaSidlinskiy I need help doing that.

OpenStudy (welshfella):

first take x out of the quadratic function

OpenStudy (inowalst):

But one is negative and one isn't..

OpenStudy (welshfella):

oh sorry you can simplify first (x^2+5x−14x^2 = 5x - 13x^2 = x(5 - 13x)

OpenStudy (welshfella):

there are no common factors here are you sure you have the question right?

OpenStudy (inowalst):

The 14 isn't supposed to have \(x^2\)

OpenStudy (welshfella):

i suspect thet the quadratic is x^2+5x−14

OpenStudy (welshfella):

right

OpenStudy (inowalst):

Sorry about that.

OpenStudy (welshfella):

to factor x^2+5x−14 you need to find 2 numbers whose product is -14 and whose sum is +5

OpenStudy (welshfella):

can you tell me what those 2 numbers sre?

OpenStudy (welshfella):

think of the factors of -14

OpenStudy (inowalst):

Their product has to be -14 but the sum has to be 5?

OpenStudy (welshfella):

yes

OpenStudy (inowalst):

Hm.. I don't really know, but I know for the -14..

OpenStudy (welshfella):

-14 = -7 * 2 or -14 = -2 * 7

OpenStudy (welshfella):

which 2 factors add up to +5?

OpenStudy (inowalst):

-2+7

OpenStudy (welshfella):

right on so the factors of (x^2+5x−14) are ( x - 7)(x + 2)

OpenStudy (welshfella):

if you expand this function you'll see it works out to x^2 + 5x - 14

OpenStudy (welshfella):

* oops my bad its (x + 7)(x - 2)

OpenStudy (inowalst):

Yes. Thank you so much.

OpenStudy (welshfella):

Now you need to factor x^2 - 49

OpenStudy (inowalst):

Okay.

OpenStudy (welshfella):

can u do that?

OpenStudy (welshfella):

49 - x^2

OpenStudy (inowalst):

(x-7)(x+7) ? I don't know..

OpenStudy (welshfella):

49 - x^2 = (7 - x)(7 + x) so can you see the common factor?

OpenStudy (inowalst):

Yes.

OpenStudy (welshfella):

x + 7

OpenStudy (inowalst):

Oh, okay. Thanks a bunch.